THE LOW-LYING BENDING VIBRONIC BANDS OF THE MgNC Ã^{2}Π – X̃ ^{2}Σ^{+} TRANSITION: ANALYSIS OF THE P-TYPE DOUBLING

\underline{MASARU FUKUSHIMA} and TAKASHI ISHIWATA, Faculty of Information Sciences, Hiroshima City University, Asa-Minami, Hiroshima 731-3194, Japan.

We have generated MgNC in supersonic free jet expansions, and measured the laser induced fluorescence excitation spectra of the Mg-N-C bending vibronic bands of the Ã ^{2}Π – X̃ ^{2}Σ^{+} transition. In addition to the two vibronic bands, 2_{0}^{1}, κ 2Σ^{(+)} – and 2_{0}^{2}, κ ^{2}Π – ^{2}Σ^{+}, reported by Wright and Miller^{a}, the 2_{0}^{2}, μ ^{2}Π_{\frac{1}{2}} – ^{2}Σ^{+} sub-band were found just above the 2_{0}^{1} vibronic band. The most remarkable feature of this newly observed sub-band is unexpected rotational structure on the upper vibronic level, Ã (020) μ ^{2}Π_{\frac{1}{2}} ; i.e. the rotational levels show splitting into the e and f sub-levels and have different rotational constants on the each rotational sub-level. Introducing the Coriolis coupling terms of −J^{±}G^{∓} and S^{±}G^{∓}, which could be called the G-uncoupling operator and the spin-Coriolis interaction, respectively, into the rotational Hamiltonian, this unexpected rotational structure has been analyzed quite reasonably.

aR. R. Wright and T. A. Miller, J. Mol. Spectrosc. \underline{194}, 219 (1999).